200 units of a product cost #800 and 400 units cost #1200. Find the fixed cost and the cost of producing 800 units.
1
Expert's answer
2020-04-23T11:54:08-0400

Let's fixed cost be x, and variable cost be y. Now we can write two equations:

"x+200y=400"

"x+400y=1200"

By solving that equations we find x=400, y=2.

So, fixed cost is 400.

The cost of producing 800 units is:

"400+2\times800=2000."